OpenStudy (anonymous):

Prove the following about greatest common divisors. gcd(a,b) denotes the greatest common divisor of a and b. i) If gcd(a,b)=1 then gcd(ac,b)=gcd(b,c) ii) If gcd(a,b)=1 and c|(a+b) then gcd(a,c)=gcd(b,c)=1

OpenStudy (anonymous):

c|(a+b) denotes that c divides (a+b), that is, there exists an integer k such that (a+b)=k.c

OpenStudy (anonymous):

what you need for both of these i think, is that if \((a,b)=1\) there exists \(x,y\) with \(ax+by=1\)

OpenStudy (anonymous):

lets call \((ac,b)=d\) you want to show that \(d|c\) since \(d|b\) and \(d|ac\) it must also divide \(acx+bcy\) for the same \(x,y\) above

OpenStudy (anonymous):

therefore \(d|acx+bcy=c(ax+by)=c\) since \(ax+by=1\) showing you that \(d|c\)

OpenStudy (anonymous):

Why do I want to show that d|c?

OpenStudy (anonymous):

I know that, d|c and d|b would mean that... Oh I see, it would divide any linear combination of b and c. But how do I guarantee that it is the smallest number that divides that linear combination?

OpenStudy (anonymous):

you show that the gcd's are equal by showing that the gcd of one pair divides the other pair, and that the gcd of the second pair divides the first pair

OpenStudy (anonymous):

But you only showed that d|c.... There's something I'm not seeing here.

OpenStudy (anonymous):

ok let me try again because all these symbols are confusing

OpenStudy (anonymous):

you want to show gcd(ac,b)=gcd(b,c)

OpenStudy (anonymous):

so if you call \((b,c)=d\) you know that \(d|b, d|c\) and so clearly \(d|ac\) and thus \(d|gcd(ac,b)\)

OpenStudy (anonymous):

in other words, it is fairly clear that the gcd of \(c\) and \(b\) divides the gcd of \(ac\) and \(b\) right?

OpenStudy (anonymous):

if not let me know

OpenStudy (anonymous):

It's because d|ac and d|b, right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so if it divides them both, it must divide their greatest common divisor, on account of the greatest common divisor is, well, the greatest

OpenStudy (anonymous):

so the hard part is to show that the gcd of ac and b also divides the gcd of c and b

OpenStudy (anonymous):

if you can show that, they must be equal i hope the logic of that is clear i have two numbers, \(d_1\) and \(d_2\) and they each divide the other then they must be the same number

OpenStudy (anonymous):

Yes, I understand that. So, now we have to prove that the gcd(ac,b)|gcd(b,c)

OpenStudy (anonymous):

yes that is what you have to do which amounts to showing that it divides c, because it obviously divides b

OpenStudy (anonymous):

And we do so by stating that, since gcd(a,b)=1, it means that \[1=ax+by\] for some x,y integers

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ah, yes, clearly divides b.

OpenStudy (anonymous):

So we multiply by c

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

So \[c=acx+bcy\], which can be rewritten as \[c=c(ax)+b(cy)\] which means that c is a multiple of the gcd(b,c), obviously. That means that gcd(b,c)|c

OpenStudy (anonymous):

hmm i think the logic got a little lost there

OpenStudy (anonymous):

Yes, i'm now stating obvious things.

OpenStudy (anonymous):

Let me get a piece of paper.

OpenStudy (anonymous):

only one typo

OpenStudy (anonymous):

since gcd (ac, b) divides ac it must divide acx likewise it must divide by therefore it must divide \(cax+by=c\)

OpenStudy (anonymous):

Why is \[cax+by=c\] ?

OpenStudy (anonymous):

because i made a typo

OpenStudy (anonymous):

should be \(cax+cby=c\)

OpenStudy (anonymous):

and you can see that gcd(ac,b) must divide both terms on the left, so it must divide c

OpenStudy (anonymous):

hope the logic is clear, sorry for the typo above

OpenStudy (anonymous):

Ok, let me write this nicely and see if I'm doing it right.